OpenStudy (jkasdhk):

If α and β are complex cube roots of unity then (α^4 β^4)+1/αβ=?

OpenStudy (anonymous):

is it 2

OpenStudy (jkasdhk):

How?

OpenStudy (anonymous):

the product of the roots αβ=1 and (α^3 β^3)=1

OpenStudy (jkasdhk):

How is (α^3 β^3)=1?

OpenStudy (jkasdhk):

We see sum of roots and product of roots?

OpenStudy (anonymous):

\[\alpha * \beta \] isnt equal to 1....cube root remember?

OpenStudy (jkasdhk):

the answer given is 0. i dont know how?

OpenStudy (anonymous):

infact the answer is zero

OpenStudy (jkasdhk):

How????

OpenStudy (anonymous):

well its complex cube root, \[\alpha = \beta \] = \[\sqrt{-1}\]. so that would make cube of it equal to 1. but the other term -1. 1-1=0....

OpenStudy (jkasdhk):

complex cube roots means α=β?

OpenStudy (anonymous):

do u agree the roots \[\alpha =\Omega and \beta = \Omega ^{2}\]? or vice versa

OpenStudy (anonymous):

no, not necessarily all the time. for unity yes.

OpenStudy (anonymous):

if α and β are the complex cube roots of 1 then α and β must satisfy the eq. x^2 +x +1=0 and x^3=1 thus α + β=-1 and αβ=1

OpenStudy (anonymous):

\[ \alpha=e^{\frac{2 i \pi }{3}}; \quad \alpha^3 =1\\ \beta =1; \quad \beta^3 =1\\ \alpha \beta \ne 1 \]

OpenStudy (jkasdhk):

@amrit110: can u please tell me the steps to solve this ques?

OpenStudy (anonymous):

\[\Omega ^{3} = 1\]

OpenStudy (anonymous):

ok. see complex cube roots of unity so \[x^{3}=1\]. so when u solve for x it is obviously the root of (-1). now this means that for this case we can safely assume alpha n beta to be equal to x. although this is not true for all cubic equations, sometimes the roots can be complex conjugate too. but it doesnt matter.

OpenStudy (shubhamsrg):

note that the complex cube roots of unity are the zeroes of the eqn x^2 + x +1 so (alpha)(beta) = 1 try now..

OpenStudy (anonymous):

There are 3 complex roots of unity: \[ \left\{1,e^{\frac{2 i \pi }{3}},e^{-\frac{2 i \pi }{3}}\right\} \]

OpenStudy (anonymous):

A complex root of unity is a root of \[ x^3 =1 \]

OpenStudy (anonymous):

@eliassaab i didn't get how come αβ not 1

OpenStudy (jkasdhk):

@amrit110: What will be the answer of (α^4 β^4) ?

OpenStudy (anonymous):

If both \( \alpha \ne 1 \) and \( \beta \ne 1 \) then \(\alpha \beta =1\)

OpenStudy (anonymous):

In the above case the ratio is 2.

OpenStudy (anonymous):

it should be 1 jkasdhk

OpenStudy (anonymous):

yes for complex nos .. for real nos If both α≠1 and β≠1 then αβ not 1

OpenStudy (jkasdhk):

and 1/αβ would be?

OpenStudy (anonymous):

1/αβ would be 1

OpenStudy (jkasdhk):

1+1 will be 2? right?

OpenStudy (jkasdhk):

but 2 is not the answer

OpenStudy (anonymous):

is the question (α^4 β^4) -1/αβ or still different

OpenStudy (anonymous):

If we consider complex roots only and \(\alpha \ne \beta\) then \[ \alpha = e^{\frac{2 i \pi }{3}}\\ \beta = e^{-\frac{2 i \pi }{3}}\\ \alpha \beta= e^0 =1 \]

OpenStudy (jkasdhk):

its (α^4 β^4) +1/αβ

OpenStudy (anonymous):

If the question is: Let \( \alpha, \beta\) be the 2 complex roots (not real), then the ratio in question is 2

OpenStudy (jkasdhk):

@amrit110: How is (α^4 β^4)=1

OpenStudy (anonymous):

\[ \alpha^4 \beta^4 = \alpha^3 \beta^3 \alpha \beta = (1)(1) =\alpha \beta =1 \] If \( \alpha\ne \beta \)

OpenStudy (anonymous):

Let \( 1, \alpha, \beta \) be the three roots of unity,then \[ \frac { \alpha ^4 \beta^4 +1}{\alpha \beta }= 2 \] It is easy to see now \[ (1) \alpha \beta =1 \] Since the last product is the product of the three roots of the polynomial \[ x^3 -1=0

OpenStudy (anonymous):

This product is \[ (-1)^3 \frac {-1}{1}=1 \] See http://en.wikipedia.org/wiki/Vieta%27s_formulas

OpenStudy (anonymous):

@jkasdhk